A 205-kg load is hung on a wire of length of 4.30 m, cross-sectional area 2.00 10-5 m2, and Young's modulus 8.00 1010 N/m2. What is its increase in length?
Answer #1

young's modulus is a ratio of tensile stress and tensile strain .
which gives , E = (F/A) /( del L/L) = FL/A 'del L'
we deduce , 'del L'= FL/AE
in ques. m=205 kg . , F = mg = 205x9.8 N =2009 N
A=area of wire = 2.00x10^-5 m2
L= 4.3 m
E = 8.00x10^10 N/m2
so , putting value we get increase in length

'del L' = FL/AE=(2009x4.3)/( 2.00x10^-5 x8.00x10^10) m

= 5.4 x 10^-3 m